# Trajectory Optimization for UAV-Based Medical Delivery with Temporal Logic Constraints and Convex Feasible Set Collision Avoidance

### v1

@@ III-D The Convex Optimization Problem
$\quad\|[v^{x}_{t},v^{y}_{t}]^{\top}\|_{2}\leq v_{\text{max}},\ \forall t\in\mathbb{Z}_{[1,T]}$
@@
This formulation consists of convex dynamics, convexified task satisfaction constraints, and locally convex approximations of non-convex environment constraints. All terms in the objective function and all constraints are either affine or convex, making the overall formulation compatible with standard convex solvers. This enables efficient and reliable computation of safe, dynamically feasible UAV trajectories that respect mission timing and spatial requirements. The proposed convex optimization algorithm for UAV trajectory planning under STL and collision avoidance constraints is summarized in Algorithm 1.
@@
1:Input: Waypoints $\{\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{K}\}$ and the associated time windows $[\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]$ for all $k\in\mathbb{Z}_{[1,K]}$. Positive weights $w_{1}$, $w_{2}$, and $w_{3}$. Initial reference trajectory $\bar{\mathbf{x}}$.
@@
IV Simulation Result
IV-A Parameter Specification
The proposed trajectory optimization framework is tested in a simulated three-dimensional urban environment. All simulations are implemented in MATLAB using the convex optimization toolbox (CVX) [20] over a discretized time horizon of $T=30$ time steps with a uniform interval $\Delta t=1.0$ seconds.
The UAV begins at the fixed initial position:
[…]
@@ IV-A Parameter Specification
TABLE I: Delivery Targets
Hospital Coordinates $(x,y,z)$ Time Window $[\tau^{\text{start}},\tau^{\text{end}}]$
$\mathbf{h}_{1}$ [17, 18, 5]17185[17,\;18,\;5][ 17 , 18 , 5 ] $[4,15]$
$\mathbf{h}_{2}$ [42, 28, 5]42285[42,\;28,\;5][ 42 , 28 , 5 ] $[10,22]$
$\mathbf{h}_{3}$ [56, 20, 5]56205[56,\;20,\;5][ 56 , 20 , 5 ] $[16,30]$
@@
TABLE II: Obstacle Specification
@@ IV-B Results and Discussions
IV-B1 Trajectory Quality and Task Completion
@@
Fig. 1 compares the optimized UAV trajectory obtained using the proposed convex approach (orange) against that generated by a standard Sequential Quadratic Programming (SQP) solver (blue) [21]. Both 3D and top-down views are shown. The convex method successfully produces a smooth, dynamically feasible trajectory that visits all three hospitals while respecting the obstacle constraints and STL-defined time windows. In contrast, the SQP-based trajectory visibly deviates near the second and third hospitals and slightly overshoots the time window boundaries, indicating a lack of robustness in satisfying temporal logic constraints.
@@
IV-B2 Fast Convergence and Optimization Efficiency
@@
IV-B3 Temporal Logic Satisfaction and Robustness
@@
IV-B4 Comparative Discussion and Practical Implications

### v2

@@ 3.3 The Convex Optimization Problem
$\|[v^{x}_{t},v^{y}_{t}]^{\top}\|_{2}\leq v_{\text{max}},\ \forall t\in\mathbb{Z}_{[1,T]}$
@@
This formula consists of convex dynamics, convexified task satisfaction constraints, and locally convex approximations of non-convex environment constraints. All terms in the objective function and all constraints are either affine or convex, making the overall formulation compatible with standard convex solvers. This enables efficient and reliable computation of safe, dynamically feasible UAV trajectories that respect mission timing and spatial requirements. The proposed convex optimization algorithm for UAV trajectory planning under STL and collision avoidance constraints is summarized in Algorithm 1.
@@
1:Input: Waypoints $\{\mathbf{h}_{1},\mathbf{h}_{2},\dots,\mathbf{h}_{K}\}$ and the associated time windows $\tau_{k}^{\text{start}},\tau_{k}^{\text{end}}]$ for all $k\in\mathbb{Z}_{[1,K]}$. Positive weights $w_{1}$, $w_{2}$, and $w_{3}$. Initial reference trajectory $\bar{\mathbf{x}}$.
@@
4 Simulation Result
4.1 Parameter Specification
The proposed trajectory optimization framework is tested in a simulated three-dimensional urban environment. All simulations are implemented in MATLAB using the convex optimization toolbox (CVX) (Grant and Boyd, 2008) over a discretized time horizon of $T=30$ time steps with a uniform interval $\Delta t=1.0$ seconds.
The UAV begins at the fixed initial position $\mathbf{x}_{1}=[0,\;0,\;5]^{\top}$, which represents a rooftop launch pad at 5 meters above ground. The UAV’s initial velocity is zero in all directions. Its motion follows the discrete-time kinematic model defined in Section 3.1, with maximum planar speed $v_{\text{max}}=5\,\text{m/s}$, and acceleration bounds $a_{\text{max}}=2\,\text{m/s}^{2}$ in all directions.
@@ 4.1 Parameter Specification
Table 1: Delivery Targets
Hospital Coordinates Time Window
$(x,y,z)$ $\tau^{\text{start}},\tau^{\text{end}}]$
$\mathbf{h}_{1}$ [17, 18, 5][17,\;18,\;5] $[4,15]$
$\mathbf{h}_{2}$ [42, 28, 5][42,\;28,\;5] $[10,22]$
$\mathbf{h}_{3}$ [56, 20, 5][56,\;20,\;5] $[16,30]$
@@
Table 2: Obstacle Specification
@@ 4.2 Results and Discussions
4.2.1 Trajectory Quality and Task Completion
Fig. 1 compares the optimized UAV trajectory obtained using the proposed convex approach (orange) against that generated by a standard Sequential Quadratic Programming (SQP) solver (blue) (Gill et al., 2005). Both 3D and top-down views are shown. The convex method successfully produces a smooth, dynamically feasible trajectory that visits all three hospitals while respecting the obstacle constraints and STL-defined time windows. In contrast, the SQP-based trajectory visibly deviates near the second and third hospitals and slightly overshoots the time window boundaries, indicating a lack of robustness in satisfying temporal logic constraints.
@@ 4.2.1 Trajectory Quality and Task Completion
4.2.2 Fast Convergence and Optimization Efficiency
@@
4.2.3 Temporal Logic Satisfaction and Robustness
@@
4.2.4 Comparative Discussion and Practical Implications